EF002B

PI: Fanti
Title: CSS/GPS Sc.

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 03:00(16/02)-19:30(16/02)
Stations: EbMcNtOnTrSmShUrWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 4

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Mantovani at Sun Feb 16 15:06:19 EST 1997) Success

Onsala ( at Sun Feb 16 16:27:50 EST 1997) Success
No known problems

Effelsberg (dgraham at Mon Feb 17 02:52:23 EST 1997) Minor failures
Formatter probably 1 sec slow. Used caltemp=10 instead of 1.4K so Tsys calculation wrong.

Westerbork (Tony Foley at Tue Feb 18 10:06:48 EST 1997) Success
Looks OK. Please note that we needed to make lots of hand edits to the array schedule to allow for gaps where sources were out of Hour Angle range.

Torun (Roman Feiler at Wed Feb 19 18:11:01 EST 1997) Minor failures
4:32-4:44, 6:14-6:34, 19:16-19:30 off source due to power failures at antenna

Jodrell (Alastair Gunn at Fri Feb 21 03:31:57 EST 1997) Failure
POLARISATIONS ARE CROSSED. Therefore no useful data.

Cambridge (Alastair Gunn at Fri Feb 21 05:20:27 EST 1997) Minor failures
No major problems. FS crashed at about 14:51 so scan 24 missed. Weather cloudy, windy. Tapes EVN00120,0AAD EVN00101,6A9D and HS004353,EE4F sent to Bonn via Woodcroft on 17/02/97.

Shanghai (Huang Xinyong at Wed Mar 5 04:45:17 EST 1997) Success
Clock offset = -0.06 msec. No problems, weather fine.

Calibration (Fredrik T Rantakyro at Thu May 8 11:26:20 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ef2b.antab

MPIfR (Scott Aaron at Tue Sep 23 14:41:56 METDST 1997) Minor failures
Shanghai 100% data loss--telescope not connected to recorder.

MPIfR (Scott Aaron at Wed Dec 17 09:49:55 MET 1997) Minor failures
Correlation complete. No fringes to Shanghai. Jodrell Bank swapped polarizations, so no data. Tracking problems on the Torun tapes, so all baselines to Torun deleted from recorrelation list. Everything else OK.